OpenStudy (anonymous):

Calculate the area of triangle ABC if the sides are 10, 14, and 15. I just dont get it??

OpenStudy (anonymous):

simple s=(10+14+15)/2=19.5 \[area=\sqrt{(19.5*9.5*5.5*4.5)}\]

OpenStudy (anonymous):

=67.7 hope its correct

OpenStudy (anonymous):

the area is \[\sqrt{s(s-a)(s-b)(s-c)} \] a,b,c are the sides

OpenStudy (anonymous):

Area of triangle = base by height over 2

OpenStudy (anonymous):

ok thank you so much :)

OpenStudy (anonymous):

1) Let a=10, b=14 and c=15 and A, B and C are the angles directly opposite the sides in triangle ABC 2) Use cosine rule to work out an angle (A) \[a ^{2}=b ^{2}+c ^{2}-2bc \cos A\] or \[\cos A=(b ^{2}+c ^{2}-a^{2})/(2bc)\] \[\cos A=(14 ^{2}+15 ^{2}-10^{2})/(2\times14\times15)\] \[\cos A=(196+225-100)/420\] \[\cos A = 0.76428....\] \[A=\cos^{-1} (0.764286...)=40.15^{o}\] 3) Area of triangle = \[(1/2)bc \sin A\] \[0.5\times14\times15\times \sin 40.15^{o}\] \[Area = 67.7\] square units